# Broadcast in Almost Mixing Time

### v2

@@ 5.2 Straightforward Lower Bounds are not Enough
Figure 1: An example where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
@@ 6 Conclusion
Our study introduces several promising avenues for future investigation. One intriguing direction involves developing distributed processes capable of rapidly covering graphs that also maintain composability, that is, efficiently supporting multiple simultaneous executions under congestion constraints. Another important direction is identifying graph properties beyond traditional metrics such as diameter and minimum cut, which would facilitate establishing tight lower bounds. Subsequently, algorithms matching these lower bounds could be devised, advancing both theoretical understanding and practical efficiency in distributed computing.
Acknowledgments
Authors are grateful to Diana Ghinea and Bernhard Haeupler for helping to improve the paper greatly.
@@ Appendix A Technical definitions
[…]
@@ Appendix B Straightforward Lower Bounds are not Enough
In this section, it will be more comfortable for us to consider a more general model than CONGEST, namely the model where edges have arbitrary bandwidth. To transform a graph with arbitrary bandwidths to a graph with all bandwidths equal to $1$, we do the following. The source $s$ corresponds to a single node in the new graph. For a node $v\neq s$ in the original graph, let $B$ denote the maximal bandwidth of its adjacent edges. In the new graph, node $v$ then corresponds to a clique of $B$ nodes. We call this clique a $v$-clique. If in the original graph nodes $v\neq s$ and $u\neq s$ were connected by an edge of bandwidth $b$, we pick (arbitrary) $b$ nodes in $v$-clique, $b$ nodes in $u$-clique and draw $b$ edges between picked nodes to establish a perfect matching. For every edge $(s,u)$ of bandwidth $b$, we connect the new source with $b$ arbitrary nodes of the $u$-clique. We call the resulting graph the corresponding CONGEST graph.
@@
The first inequality is straightforward. We prove the second inequality by induction on the length of the path, that is if there is a path in $G$ from $u$ to $v$ of length $l$, then for any nodes $u^{\prime}$ and $v^{\prime}$ in $u$-clique and $v$-clique respectively, there is a path between $u^{\prime}$ and $v^{\prime}$ in $G^{\prime}$ of length $12l+$ 1.
@@
[…]
@@
Consider the original graph $G$ and its corresponding CONGEST graph $G^{\prime}$. Together with set $M$ of messages, they define a multi-message broadcast problem in generalized CONGEST and CONGEST respectively. Let $\mathrm{OPT}(G)$ and $\mathrm{OPT}(G^{\prime})$ denote the optimal round complexities for $G$ and $G^{\prime}$ respectively. Then $\mathrm{OPT}(G)\leq\mathrm{OPT}(G^{\prime})$.
@@
Consider an execution $E^{\prime}$ for $G^{\prime}$ which achieves $\mathrm{OPT}$. We claim that we can build an execution $E$ for $G$, such that for every round $r$ of $E^{\prime}$ and for every $v\in V(G)$, after round $r$ in $E$ $v$ knows all the messages that know the nodes of $v$-clique after round $r$ in $E^{\prime}$. To do so, consider round $r$ and some $v\in V(G)$. Let us say that nodes in $v$-clique in $E^{\prime}$ in round $r$ receive messages $M_{1}$ from the $u_{1}$-clique, messages $M_{2}$ from $u_{2}$-clique and so forth for all neighboring cliques. Then, in $E$ $u_{i}$ sends $M_{i}$ to $v$ satisfying the invariant. Note that $u_{i}$ can do this in terms of the bandwidth by construction of the corresponding CONGEST graph, and in terms of knowing $M_{i}$ by the invariant.
Figure 3: An example graph $G$ where diameter and minimum cut are not telling. Here edge labels denote bandwidth.
@@
First, note that $D(G)=4$, hence by Claim 2, $D(G^{\prime})=O(1)$. Second, notice that $minCut(G)=m$ and the minimal maximal bandwidth of an edge adjacent to some node in $V(G)\setminus\{s\}$ is equal to $m$, therefore, by Claim 3, $m-1\leq minCut(G^{\prime})\leq m$. Finally, note that $\Delta(G\setminus s)=3$, hence by Claim 4, $\mathrm{OPT}(G^{\prime})\geq\mathrm{OPT}(G)$, where $\mathrm{OPT}$ is the optimal round complexity. Therefore, it is sufficient to show that $\mathrm{OPT}(G)=\Omega(\sqrt{m})$.
To see this, let us label the “bottom“ nodes of $G$ as $t_{1},t_{2},\ldots,t_{m}$, and let us focus on $t_{1}$. We claim that $\Omega(\sqrt{m})$ rounds are needed for $t_{1}$ only to get to know $M$ (become saturated). For the sake of contradiction, assume that we can saturate $t_{1}$ in $\leq\sqrt{m}-1$ rounds. That means, that it can be saturated without using the edges $(v_{\sqrt{m}+1},t_{\sqrt{m}+1})$ and $(v_{\sqrt{m}+1},t_{\sqrt{m}+2})$. But if we remove those edges, $minCut(s,t_{1})\leq\sqrt{m}$, implying that the number of rounds needed is at least $\frac{m}{\sqrt{m}}=\sqrt{m}$, a contradiction.
@@ Appendix C NP-hardness
To show the NP-hardness of a multi-message broadcast problem, we will also use a generalization of CONGEST that allows for arbitrary edge bandwidth, though this time the construction is different. In this section, we will consider a specific layered graph with layers induced by the distance from $s$. In that graph all edges connect nodes of consecutive layers. This graph has arbitrarily large bandwidths assigned to its edges, so we transform it into a graph with unit bandwidths by doing the following. For each node $v$ on layer $0<l<\text{max layer}$, we crea te a group of $n$ nodes called $v_{out}$, where $n$ denotes the number of messages (we change the notation due to reduction). Then, for every edge $(u,v)$ of the original graph, where $u$ belongs to the previous layer ($l-1$), if that edge has bandwidth $b\leq n$, we create a group of $b$ nodes called $v_{u-in}$ and we connect arbitrary $b$ nodes of $u_{out}$ 1 to 1 to node of $v_{u-in}$. For every $u$ we connect every node of $v_{u-in}$ to every node of $v_{out}$. For node $s$, we replace it with a new sink $s^{\prime}$ and create a $K_{n,n}$ with its first half called $s_{in}$ and its second half called $s_{out}$. We then connect $s^{\prime}$ to all the nodes in $s_{in}$ and we connect all the nodes of $s_{out}$ to the $in$-s of the nodes $s$ is connected to in the original graph in a way described above. For all the nodes of the last layer, we keep them a single node and draw all the incoming edges to this node. We call the resulting graph of this transformation the transformed graph. Please see Figure 4 for an example.
Figure 4: Example of transforming a layered graph with arbitrary bandwidths into the graph suitable for CONGEST. Here the number of messages $n$ is $3$.
@@
Consider a layered graph $G$ with arbitrary bandwidths and $G^{\prime}$ being its transformed version. Denote $l$ to be the depth of $G$ and let $T$ be the set of nodes in $G$ in layer $l$. Then it is possible to saturate all nodes in $T$ in $l$ rounds in $G$ if and only if it is possible to saturate all nodes in $T$ in $12l+$ 1 rounds in $G^{\prime}$.
@@
Consider an execution $E$ for $G$ in which all nodes in $T$ are saturated in $l$ rounds. We build an execution $E^{\prime}$ for $G^{\prime}$ that satisfies the following invariant: for $0\leq r<l$, after $22r+$ 2 rounds of $E^{\prime}$, for every node $v\in V(G)$ such that $v$ is in layer $t\leq r$, nodes in $v_{out}$ know the same set of messages in $E^{\prime}$ as $v$ knows in $E$ after round $r$. For $r=0$, we make $s^{\prime}$ send all messages to $s_{in}$ (a distinct message to each node) and $s_{in}$ to relay those messages to $s_{out}$. Then, if in round $r>0$ in $E$ $u$ sends $v$ $b$ messages, $u_{out}$ send $v_{u-in}$ those $b$ messages and then $v_{u-in}$ relay those to $v_{out}$. In the final $l$-th round of $E$, nodes of $G$ send messages to $t_{i}\in T$. This can be simulated in $E^{\prime}$ within one round, making it $2(l-1)+2=2l$ rounds to reach $v_{out}$ for all $v$-s in layer $l-1$ and $1$ more round to saturate $T$.
The proof of the other direction proceeds analogously maintaining the invariant that every node $v\in V(G)$ in $E$ after $r$ rounds knows all the messages that $v_{out}$ knows in $E^{\prime}$ after $22r+$ 2 rounds.
@@
With a slight abuse of notation, we introduce two other nodes, namely $S_{1}$ and $S_{2}$ that intuitively correspond to a partition of $S$. We focus on solving the set partition problem for the given size of the parts, i.e. $|S_{1}|=n_{1}$ and $|S_{2}|=n_{2}$ with $n_{1}+n_{2}=n$. Obviously, this version is also NP-complete. We draw an edge $(s,S_{1})$ of bandwidth $n_{1}$ and $(s,S_{2})$ of bandwidth $n_{2}$. We want $S_{1}$ and $S_{2}$ to be in layer $2$, so we introduce intermediate nodes on those edges whose role will simply be to relay messages.
@@
We introduce nodes $t_{i,1}$ and $t_{i,2}$ for $i\in[m]$. For $i\in[m]$ we draw an edge $(u_{i,1},t_{i,1})$ of bandwidth $n$ and, and this is the crux of the reduction, an edge $(s,t_{i,1})$ of bandwidth $n-n_{1}\mathbf{-1}$ and of length $4$ (with $3$ intermediate nodes). The idea here is that $t_{i,1}$ can be saturated after round $4$ if and only if it receives more than $n_{1}$ messages from $u_{i,1}$ implying $F[i]\not\subset S_{1}$. Similarly, we do for $S_{2}$. See the resulting construction in Figure 5.
Figure 5: Graph $G$ with arbitrary bandwidths for which it is NP-hard to optimally solve multi-message broadcast. Some nodes are depicted in gray since they serve no other purpose but to layer the graph and in reasonable executions should only relay messages.
If we now consider a transformed graph $G^{\prime}$, we want to focus on saturating nodes in $T=\{t_{11},t_{12},\ldots,t_{m1},t_{m2}\}$, though in multi-message broadcast problem the goal is to saturate all nodes. To account for that, for each node $v\in V(G^{\prime})\setminus T$, we will make sure that it can be saturated in $9$ rounds. We do that by introducing a path of length $9$ and bandwidth $n$ from $s$ to $v$. In particular, each such path has $6$ intermediate layers of $n$ nodes each. Each node in the first layer is connected to each node in $s_{out}$. Each node in layer $1<l\leq 6$ is connected to each node in layer $l-1$, and $v$ is connected to each node in layer $6$. This way we obtain graph $G^{\prime\prime}$. Note that introducing these additional paths does not help saturating $T$ in less than $9$ rounds, that is $T$ can be saturated in $9$ rounds in $G^{\prime\prime}$ iff it can be saturated in $9$ rounds in $G^{\prime}$.
This observations combined with Claim 5 allow us to establish the following sequence of equivalent statements ($\Leftrightarrow$ denotes equivalence):
@@
First, assume it is possible to split $S$ and this splitting is $S_{1}$ and $S_{2}$. Then we claim it is possible to saturate $T$ in $G$ in $4$ rounds. To do so, let $s$ send $\{m_{i}\mid i\in S_{1}\}$ to $S_{1}$ and $\{m_{i}\mid i\in S_{2}\}$ to $S_{2}$. Also, let it send $m_{i}$ to $v_{i}$ and to $t_{ij}$, $i\in[m]$, $j\in\{1,2\}$, $s$ sends $S\setminus(F_{i}\cup S_{j})$. After that, nodes only relay the messages they have to further layers. Now we claim that after round $4$, all nodes in $T$ are saturated. Indeed, for instance, $t_{i1}$ will receive $F_{i}\cup S_{1}\cup(S\setminus(F_{i}\cup S_{1}))=S$, the main point being that since $F_{i}\subsetneq S_{1}$, $|F_{i}\cup S_{1}|>|S_{1}|=n_{1}$, therefore $|S\setminus(F_{i}\cup S_{1})|\leq n-n_{1}-1$ and $s$ can send it whole.
Now, assume we can saturate $T$ in $G$ in $4$ rounds. This implies that every $u_{i1}$ in round $3$ holds more than $n_{1}$ messages, implying that messages held by $F_{i}$ are not a strict subset of messages held by $S_{1}$ in round $2$. Analogously, it holds for $F_{i}$ and $S_{2}$. This means, there is (possibly non-injective) mapping $\phi$ of $\{v_{1},\ldots,v_{n}\}$ into $S$ so that $\forall i\in[m],j\in\{1,2\}$ it holds that ()($\ast$)( ∗ ) $\bigcup\limits_{l\in F[i]}\phi(v_{l})\subsetneq S_{j}$. Note that by making $\phi$ injective (and thus bijective) by iteratively taking a colliding pair $x,y$ ($\phi(x)=\phi(y)$) and assigning $y$ to the so far uncovered element, we cannot break $\ast$. Therefore, we can assume that $\phi$ (i.e., distribution of messages across $v_{i}$) is bijective, which gives a solution to the splitting problem up to permuting the elements.

### v3

@@ 5.2 Straightforward Lower Bounds are not Enough
Figure 1: An example where diameter and minimum cut are not telling. Edge labels denote bandwidth.
@@ 6 Conclusion
Our study introduces several promising avenues for future investigation. One intriguing direction involves developing distributed processes capable of rapidly covering graphs that also maintain composability, that is, efficiently supporting multiple simultaneous executions under congestion constraints. Another important direction is identifying graph properties beyond traditional metrics such as diameter and minimum cut, which would facilitate establishing tight lower bounds. Subsequently, algorithms matching these lower bounds could be devised, closing the open problem of multi-message broadcast.
@@ Appendix A Technical definitions
$\Pr\left(X\leq(1-\delta)\mu\right)$ $\leq e^{-\delta^{2}\mu/2},\ 0\leq\delta$
$\Pr\left(X\geq(1+\delta)\mu\right)$ $\leq e^{-\delta^{2}\mu/(2+\delta)},\ 0\leq\delta$
$\Pr\left(|X-\mu|\geq\delta\mu\right)$ $\leq 2e^{-\delta^{2}\mu/3},\ 0\leq\delta\leq 1.$
@@ Appendix B Straightforward Lower Bounds are not Enough
In this section, it will be more comfortable for us to consider a more general model than CONGEST, namely the model where edges have arbitrary bandwidth. To transform a graph with arbitrary bandwidths to a graph with all bandwidths equal to $1$, we do the following. The source $s$ corresponds to a single node in the new graph. For a node $v\neq s$ in the original graph, let $B$ denote the maximal bandwidth of its adjacent edges. In the new graph, node $v$ then corresponds to a clique of $B$ nodes. We call this clique a $v$-clique. If in the original graph nodes $v\neq s$ and $u\neq s$ were connected by an edge of bandwidth $b$, we pick (arbitrary) $b$ nodes in $v$-clique, $b$ nodes in $u$-clique, and draw $b$ edges between picked nodes to establish a perfect matching. For every edge $(s,u)$ of bandwidth $b$, we connect the new source with $b$ arbitrary nodes of the $u$-clique. We call the resulting graph the corresponding CONGEST graph.
@@
The first inequality is straightforward. We prove the second inequality by induction on the length of the path, that is if there is a path in $G$ from $u$ to $v$ of length $l$, then for any nodes $u^{\prime}$ and $v^{\prime}$ in $u$-clique and $v$-clique respectively, there is a path between $u^{\prime}$ and $v^{\prime}$ in $G^{\prime}$ of length $2l+1$.
@@
The second inequality is straightforward. For the first inequality, note that each cut of $G^{\prime}$ either cuts some clique or does not. In case it does not, it corresponds to a cut in $G$ and has the same size. In case it does, it is at least the size of the induced cut for that clique, which is at least $\min\limits_{v\in V(G)\setminus\{s\}}\text{size of the $v$-clique}-1$.
@@
Consider the original graph $G$ and its corresponding CONGEST graph $G^{\prime}$. Together with a set $M$ of messages, they define a multi-message broadcast problem in generalized CONGEST and CONGEST, respectively. Let $\mathrm{OPT}(G)$ and $\mathrm{OPT}(G^{\prime})$ denote the optimal round complexities for $G$ and $G^{\prime}$ respectively. Then $\mathrm{OPT}(G)\leq\mathrm{OPT}(G^{\prime})$.
@@
Consider an execution $E^{\prime}$ for $G^{\prime}$ which achieves $\mathrm{OPT}$. We claim that we can build an execution $E$ for $G$, such that for every round $r$ of $E^{\prime}$ and for every $v\in V(G)$, after round $r$ in $E$, $v$ knows all the messages that the nodes of $v$-clique know after round $r$ in $E^{\prime}$. To do so, consider a round $r$ and some $v\in V(G)$. Let us say that nodes in $v$-clique in $E^{\prime}$ in round $r$ receive messages $M_{1}$ from the $u_{1}$-clique, messages $M_{2}$ from the $u_{2}$-clique, and so forth for all neighboring cliques. Then, in $E$ $u_{i}$ sends $M_{i}$ to $v$ satisfying the invariant. Note that $u_{i}$ can do this in terms of the bandwidth by construction of the corresponding CONGEST graph, and in terms of knowing $M_{i}$ by the invariant.
@@
First, note that $D(G)=2$, hence by Claim 2, $D(G^{\prime})=O(1)$. Second, notice that $minCut(G)=k$ and the minimal maximal bandwidth of an edge adjacent to some node in $V(G)\setminus\{s\}$ is equal to $k$, therefore, by Claim 3, $k-1\leq minCut(G^{\prime})\leq k$. Finally, by Claim 4, $\mathrm{OPT}(G^{\prime})\geq\mathrm{OPT}(G)$, where $\mathrm{OPT}$ is the optimal round complexity. Therefore, it is sufficient to show that $\mathrm{OPT}(G)=\Omega(\sqrt{k})$.
Figure 3: An example graph $G$ where diameter and minimum cut are not telling. Here, edge labels denote bandwidth.
We claim that $\Omega(\sqrt{k})$ rounds are needed for $v_{1}$ only to get to know $M$ (become saturated). For the sake of contradiction, assume that we can saturate $v_{1}$ in $\leq\sqrt{k}-1$ rounds. That means that it can be saturated without using the edges $(s,v_{\sqrt{k}+1})$ and $(v_{\sqrt{k}+1},v_{\sqrt{k}+2})$. But if we remove those edges, $minCut(s,v_{1})\leq\sqrt{k}$, implying that the number of rounds needed is at least $\frac{k}{\sqrt{k}}=\sqrt{k}$, a contradiction.
@@ Appendix C NP-hardness
To show the NP-hardness of a multi-message broadcast problem, we will also use a generalization of CONGEST that allows for arbitrary edge bandwidth, though this time the construction is different. In this section, we will consider a specific layered graph with layers induced by the distance from $s$. In that graph, all edges connect nodes of consecutive layers. This graph has arbitrarily large bandwidths assigned to its edges, so we transform it into a graph with unit bandwidths by doing the following. For each node $v$ on layer $0<l<\text{max layer}$, we create a group of $n$ nodes called $v_{out}$, where $n$ denotes the number of messages (we change the notation due to reduction). Then, for every edge $(u,v)$ of the original graph, where $u$ belongs to the previous layer (l1l-1), if that edge has bandwidth $b\leq n$, we create a group of $b$ nodes called $v_{u-in}$ and we connect arbitrary $b$ nodes of $u_{out}$ 1 to 1 to node of $v_{u-in}$. For every $u$, we connect every node of $v_{u-in}$ to every node of $v_{out}$. For node $s$, we replace it with a new sink $s^{\prime}$ and create a $K_{n,n}$ with its first half called $s_{in}$ and its second half called $s_{out}$. We then connect $s^{\prime}$ to all the nodes in $s_{in}$, and we connect all the nodes of $s_{out}$ to the $in$-s of the nodes $s$ is connected to in the original graph in a way described above. For all the nodes of the last layer, we keep them a single node and draw all the incoming edges to this node. We call the resulting graph of this transformation the transformed graph. Please see Figure 4 for an example.
Figure 4: Example of transforming a layered graph with arbitrary bandwidths into a graph suitable for CONGEST. Here the number of messages $n$ is $3$.
@@
Consider a layered graph $G$ with arbitrary bandwidths and $G^{\prime}$ being its transformed version. Denote $l$ to be the depth of $G$ and let $T$ be the set of nodes in $G$ in layer $l$. Then it is possible to saturate all nodes in $T$ in $l$ rounds in $G$ if and only if it is possible to saturate all nodes in $T$ in $2l+1$ rounds in $G^{\prime}$.
@@
Consider an execution $E$ for $G$ in which all nodes in $T$ are saturated in $l$ rounds. We build an execution $E^{\prime}$ for $G^{\prime}$ that satisfies the following invariant: for $0\leq r<l$, after $2r+2$ rounds of $E^{\prime}$, for every node $v\in V(G)$ such that $v$ is in layer $t\leq r$, nodes in $v_{out}$ know the same set of messages in $E^{\prime}$ as $v$ knows in $E$ after round $r$. For $r=0$, we make $s^{\prime}$ send all messages to $s_{in}$ (a distinct message to each node) and $s_{in}$ to relay those messages to $s_{out}$. Then, if in round $r>0$ in $E$ $u$ sends $v$ $b$ messages, $u_{out}$ send $v_{u-in}$ those $b$ messages and then $v_{u-in}$ relay those to $v_{out}$. In the final $l$-th round of $E$, nodes of $G$ send messages to $t_{i}\in T$. This can be simulated in $E^{\prime}$ within one round, making it 2(l1)+2=2l2(l-1)+2=2l rounds to reach $v_{out}$ for all $v$-s in layer l1l-1 and $1$ more round to saturate $T$.
The proof of the other direction proceeds analogously, maintaining the invariant that every node $v\in V(G)$ in $E$ after $r$ rounds knows all the messages that $v_{out}$ knows in $E^{\prime}$ after $2r+2$ rounds.
@@
With a slight abuse of notation, we introduce two other nodes, namely $S_{1}$ and $S_{2}$, that intuitively correspond to a partition of $S$. We focus on solving the set partition problem for the given size of the parts, i.e., $|S_{1}|=n_{1}$ and $|S_{2}|=n_{2}$ with $n_{1}+n_{2}=n$. Obviously, this version is also NP-complete. We draw an edge $(s,S_{1})$ of bandwidth $n_{1}$ and $(s,S_{2})$ of bandwidth $n_{2}$. We want $S_{1}$ and $S_{2}$ to be in layer $2$, so we introduce intermediate nodes on those edges whose role will simply be to relay messages.
@@
We introduce nodes $t_{i,1}$ and $t_{i,2}$ for $i\in[m]$. For $i\in[m]$ we draw an edge $(u_{i,1},t_{i,1})$ of bandwidth $n$ and, and this is the crux of the reduction, an edge $(s,t_{i,1})$ of bandwidth $n-n_{1}\mathbf{-1}$ and of length $4$ (with $3$ intermediate nodes). The idea here is that $t_{i,1}$ can be saturated after round $4$ if and only if it receives more than $n_{1}$ messages from $u_{i,1},$ implying $F[i]\not\subset S_{1}$. Similarly, we do for $S_{2}$. See the resulting construction in Figure 5.
Figure 5: Graph $G$ with arbitrary bandwidths for which it is NP-hard to optimally solve multi-message broadcast. Some nodes are depicted in gray since they serve no other purpose but to layer the graph, and in reasonable executions, should only relay messages.
If we now consider a transformed graph $G^{\prime}$, we want to focus on saturating nodes in $T=\{t_{11},t_{12},\ldots,t_{m1},t_{m2}\}$, though in multi-message broadcast problem the goal is to saturate all nodes. To account for that, for each node $v\in V(G^{\prime})\setminus T$, we will make sure that it can be saturated in $9$ rounds. We do that by introducing a path of length $9$ and bandwidth $n$ from $s$ to $v$. In particular, each such path has $6$ intermediate layers of $n$ nodes each. Each node in the first layer is connected to each node in $s_{out}$. Each node in layer $1<l\leq 6$ is connected to each node in layer l1l-1, and $v$ is connected to each node in layer $6$. This way, we obtain the graph $G^{\prime\prime}$. Note that introducing these additional paths does not help saturate $T$ in fewer than $9$ rounds, that is, $T$ can be saturated in $9$ rounds in $G^{\prime\prime}$ iff it can be saturated in $9$ rounds in $G^{\prime}$.
These observations, combined with Claim 5 allow us to establish the following sequence of equivalent statements ($\Leftrightarrow$ denotes equivalence):
@@
First, assume it is possible to split $S$, and this splitting is $S_{1}$ and $S_{2}$. Then we claim it is possible to saturate $T$ in $G$ in $4$ rounds. To do so, let $s$ send $\{m_{i}\mid i\in S_{1}\}$ to $S_{1}$ and $\{m_{i}\mid i\in S_{2}\}$ to $S_{2}$. Also, let it send $m_{i}$ to $v_{i}$ and to $t_{ij}$, $i\in[m]$, $j\in\{1,2\}$, $s$ sends $S\setminus(F_{i}\cup S_{j})$. After that, nodes only relay the messages they have to further layers. Now we claim that after round $4$, all nodes in $T$ are saturated. Indeed, for instance, $t_{i1}$ will receive $F_{i}\cup S_{1}\cup(S\setminus(F_{i}\cup S_{1}))=S$, the main point being that since $F_{i}\subsetneq S_{1}$, $|F_{i}\cup S_{1}|>|S_{1}|=n_{1}$, therefore $|S\setminus(F_{i}\cup S_{1})|\leq n-n_{1}-1$ and $s$ can send it whole.
Now, assume we can saturate $T$ in $G$ in $4$ rounds. This implies that every $u_{i1}$ in round $3$ holds more than $n_{1}$ messages, implying that messages held by $F_{i}$ are not a strict subset of messages held by $S_{1}$ in round $2$. Analogously, it holds for $F_{i}$ and $S_{2}$. This means, there is (possibly non-injective) mapping $\phi$ of $\{v_{1},\ldots,v_{n}\}$ into $S$ so that $\forall i\in[m],j\in\{1,2\}$ it holds that $(\ast)$ $\bigcup\limits_{l\in F[i]}\phi(v_{l})\subsetneq S_{j}$. Note that by making $\phi$ injective (and thus bijective) by iteratively taking a colliding pair $x,y$ ($\phi(x)=\phi(y)$) and assigning $y$ to the so far uncovered element, we cannot break $\ast$. Therefore, we can assume that $\phi$ (i.e., distribution of messages across $v_{i}$) is bijective, which gives a solution to the splitting problem up to permuting the elements.